According to Dr. Aaron Flosom of the University of Minnesota your maximum ideal waist size is directly proportional to your hip size. For a woman with 40 inch hips, the maximum ideal waist size is 32 inches. What is the maximum ideal waist size for a woman with 35 inch hips.
Answer by checkley77(12844)

40/32=35/x
40x=32*35
40x=1,120
x=1,120/40
x=29 inches for the waist.